# Coward: Collision-based Watermark for Proactive Federated Backdoor Detection

### v3

@@ 4.2 Watermark Interaction and Detection
Watermark Interaction. After the watermark is injected into the global model $\bm{w}^{t}$, it is delivered to all participated clients for local training. Benign clients perform standard training on their local datasets, optimizing only for task loss. In contrast, malicious clients additionally optimize backdoor poisoning loss. Specifically, we have:
@@
where $\mathcal{D}_{k}$ denotes the $k$-th client’s clean data and $\mathcal{D}_{k}^{tri}$ is the triggered subset with label $y_{a}$ as the attacker-specified target label. During training, benign clients would only forget the OOD watermark slightly while malicious clients tend to suppress the watermark more aggressively, as their injected backdoor interfere with the watermark. This divergence in watermark retention behavior serves as the basis for our subsequent backdoor client detection. We provide a detailed analysis of these effects in Appendix  A.
Watermark Detection. After receiving trained local models from clients, the server evaluates each model’s response to the watermark task by measuring the proportion of watermarked samples predicted as the designated label $y_{m}$. The BatchNorm statistics are also switched from those used for the FL main task to those corresponding to the watermark planting phase during inspection. After collecting the watermark accuracy of all clients, a threshold $\beta$ is applied to distinguish between benign and malicious clients. Thus, clients whose watermark accuracy falls below $\beta$ are excluded from the aggregation process. Thus we get the model aggregation operation as:
@@
where $\mathbb{I}$ denotes the indicator function, and ACC represents the corresponding watermark accuracy.
@@ 5.1 Main Settings
Datasets and Models. Image classification has served as a standard benchmark task for evaluating multimedia adversarial and forensic mechanisms. Following common practice in prior work, we conduct all experiments on three image benchmark datasets, EMNIST [cohen2017emnist], CIFAR-10, and CIFAR-100 [krizhevsky2009learning], using the ResNet-18 [he2016deep] architecture.
@@
Federated Training Configurations. All experiments adopt the standard FedAvg setting [mcmahan2017fedavg], where we simulate 100 clients in total, with 10 randomly selected to participate in each training round. To simulate the inherent data heterogeneity across clients, we follow [li2020fedprox] to use Dirichlet sampling to partition the dataset. The Dirichlet concentration parameter $\alpha$ controls the label distribution skewness, where smaller values indicate more unbalanced distributions. Specifically, we use $\alpha=0.9$ by default to simulate mild skewness, and decrease it to 0.3 to model more severe distribution shifts. During local training, each benign client performs 2 epochs with a learning rate of 0.03, while the global training proceeds for 1200 rounds. For conciseness, we omit the less critical parameter details and follow [li2024backdoorindicator] for all settings not explicitly specified.
@@
Backdoor Attack Configurations. To ensure a comprehensive evaluation, we consider both attack scenarios of single-attacker and multi-attacker, each covering various attack strategies and backdoor types. (1) In the single-attacker scenario, we adopt a range of attack baselines, including the naive Vanilla training [gu2020badnets], the stealthy PGD attack [zitengsun01pgd], and the latest state-of-the-art methods Neurotoxin [Zhengming01Neurotoxin] and Chameleon [chameleon]. Regarding the backdoor type, we adopt three diverse types of triggers, ranging from the traditional visible trigger (BadNet [gu2020badnets]), to the invisible trigger (Blend [chen2017targeted]), and the natural feature-based trigger (Semantic [bagdasaryan01]), with increasing levels of imperceptibility. (2) In the multi-attacker scenario, we increase the ratio of attackers from 30% to 70%. To ensure comprehensive coverage, we include both aligned and divergent attack goals: Uniform [foolsgold] with homogeneous attackers, DBA [xie2019dba] with a shared target but distributed triggers, and NBA [nguyen2024nba] with non-cooperative attackers using distinct backdoors. Across all settings, poisoning uses a consistent learning rate of 0.03 by default, aligned with benign training, and reduces it to half for more challenging stealthy evaluation purposes. The backdoor target is class 0 by default. Similar to BackdoorIndicator, the malicious client launches the attack at the 1000th round, with poisoning action lasting for 200 rounds.
@@
Backdoor Detection Configurations. We evaluate six representative federated backdoor detection methods that span a range of strategies, from classical model-space analysis to more advanced and proactive techniques: We begin with traditional methods that rely on the statistical properties of model updates: (1) MultiKrum [multiKrum] detects backdoor models based on the Euclidean distance among client updates. (2) Foolsgold [foolsgold] identifies malicious clients by analyzing pairwise cosine similarity to detect overly aligned update directions. (3) Rflbat [wang2022rflbat] projects model updates into a PCA space and flags outliers as potential attackers. Then we further include more advanced approaches that incorporate prediction behavior or mitigation mechanisms: (4) Deepsight [rieger2022deepsight] evaluates similarity in model predictions and output neuron activations to reveal abnormal client behavior. (5) Flame [nguyen2022flame] combines detection based on model similarity together with mitigation techniques such as gradient clipping and noise injection. Finally, we include the pioneering proactive method that shifts the detection paradigm: (6) BackdoorIndicator [li2024backdoorindicator] injects OOD patterns into the global model and identifies backdoor models based on their reactions to these patterns.
@@
Configuration of Coward. Here we detail the key factors related to watermark injection and detection. To construct the OOD planting set, we use EMNIST (expanded to three channels) when the main task is CIFAR-10 or CIFAR-100, and grayscale CIFAR-10 when the main task is EMNIST. The same planting set is used for Indicator to ensure a fair comparison. By default, the planting set contains 1,000 samples. For watermark injection, we use 20% planting set samples for trigger-mapping, with a learning rate of 0.001 for 5 iterations. In our main experiments, we adopt WaNet [nguyen2021wanet] as the default watermark trigger. For detection, a fixed 5% threshold is used across all settings: clients with watermark accuracy below this are identified as malicious. The watermark target is set to 8 by default.
@@ 5.2 Main Results
For a comprehensive evaluation, We evaluate all baseline detection methods from following three perspectives:
@@
TABLE II: Detection performance against vanilla single-client backdoor attack under different non-i.i.d. settings. Boldface indicates the best result value (excluding the collapsed case with TPR/FPR=0/0), while red highlights metric-specific failures based on a 50% threshold ($i.e.$, values above 50% for FPR and ASR, and below 50% for TPR and BA).
@@
Dataset Non-i.i.d. MultiKrum FoolsGold Rflbat DeepSight Flame Indicator Coward (Ours)
@@
CIFAR-10 0.9 93.5 34.1 34.1 92.1 44.0 44.3 91.3 92.1 2.5 9.6 91.3 91.8 74.0 60.3 58.2 91.9 99.0 33.4 9.9 92.1 97.5 23.0 9.4 91.2 99.0 3.9 10.3 90.5
0.6 62.0 37.6 88.9 91.9 47.5 50.3 89.5 91.4 1.5 10.4 91.3 92.0 66.0 53.2 66.4 92.0 69.0 36.8 87.2 91.8 94.0 26.1 35.3 91.5 99.0 9.2 9.9 90.8
0.3 57.5 38.1 90.2 91.7 60.5 57.2 84.5 90.8 1.0 9.2 91.2 91.9 57.0 44.8 72.7 91.8 62.0 37.5 88.5 91.8 93.0 31.0 24.2 91.5 99.0 9.6 10.0 90.4
CIFAR-100 0.9 98.0 22.4 1.1 69.7 0.0 0.0 82.9 69.9 12.0 22.6 83.2 69.7 39.0 32.6 46.8 69.8 97.0 22.6 1.1 69.7 95.0 47.0 1.3 69.0 98.0 2.9 1.0 69.1
0.6 96.0 22.7 1.0 69.9 0.0 0.0 82.7 69.5 11.0 22.4 83.4 69.6 55.0 47.8 41.5 69.5 95.0 22.8 1.0 69.8 81.0 47.8 56.1 69.0 100.0 4.3 1.0 69.2
0.3 2.0 33.1 82.0 69.5 0.0 0.0 81.5 69.3 6.0 23.3 82.4 69.2 75.0 65.3 25.2 69.7 5.0 32.8 81.6 69.3 76.0 43.6 46.8 67.8 98.0 8.6 1.0 68.9
EMNIST 0.9 100.0 22.2 10.0 99.8 52.0 39.1 99.9 99.7 16.0 22.7 100.0 99.7 57.0 45.9 27.4 99.7 100.0 22.0 10.0 99.7 99.0 57.8 10.0 99.6 98.0 1.6 10.0 99.7
0.6 98.0 22.4 10.0 99.8 35.0 41.7 99.7 99.7 11.0 23.1 99.8 99.7 56.0 49.4 47.8 99.7 97.0 22.3 10.0 99.7 96.0 52.7 10.8 99.6 98.0 1.6 10.0 99.7
0.3 9.0 32.3 100.0 99.7 40.0 48.2 99.5 99.8 1.0 25.1 99.9 99.7 49.0 47.0 47.3 99.8 9.0 32.1 100.0 99.7 96.0 45.6 10.2 99.6 98.0 8.3 10.0 99.7
TABLE III: Detection performance against stealthy single-client backdoor attacks with advanced training strategies. Result highlighting (in boldface and in red) follows the same rules as those described in Table II.
Dataset Attack MultiKrum FoolsGold Rflbat DeepSight Flame Indicator Coward (Ours)

### v4

@@ 4.2 Watermark Interaction and Detection
Watermark Interaction. After the watermark is injected into the global model $\bm{w}^{t}$, it is delivered to all participating clients for local training. Benign clients perform standard training on their local datasets, optimizing only for task loss. In contrast, malicious clients additionally optimize backdoor poisoning loss. Specifically, we have:
@@
where $\mathcal{D}_{k}$ denotes the $k$-th client’s clean data and $\mathcal{D}_{k}^{tri}$ is the triggered subset with label $y_{a}$ as the attacker’s target label. During training, benign clients only slightly forget the OOD watermark while malicious clients tend to suppress the watermark more aggressively, as their injected backdoors interfere with the watermark. This divergence in watermark retention behavior serves as the basis for our subsequent backdoor client detection. We provide a detailed analysis of these effects in Supplementary Section A.
Watermark Detection. After receiving trained local models from clients, the server evaluates each model’s response to the watermark task by measuring the proportion of watermarked samples predicted as the designated label $y_{m}$. The BatchNorm statistics are also switched from those used for the FL main task to those corresponding to the watermark planting phase during inspection. After collecting the watermark accuracy of all clients, a threshold $\beta$ is applied to distinguish between benign and malicious clients. Thus, clients whose watermark accuracy falls below $\beta$ are excluded from the aggregation process. Thus, we obtain the model aggregation operation as:
@@
where $\mathbb{I}$ denotes the indicator function, and ACC represents the watermark accuracy.
@@ 5.1 Main Settings
Datasets and Models. Image classification has served as a standard benchmark task for evaluating multimedia adversarial and forensic mechanisms. Following common practice in prior work, we conduct all experiments on three image benchmark datasets, EMNIST, CIFAR-10, and CIFAR-100, using the ResNet-18 architecture.
@@
Federated Training Configurations. All experiments adopt the standard FedAvg setting [25], where we simulate 100 clients in total, with 10 randomly selected to participate in each training round. To simulate the inherent data heterogeneity across clients, we follow [18] to use Dirichlet sampling to partition the dataset. The Dirichlet concentration parameter $\alpha$ controls the label distribution skewness, where smaller values indicate more unbalanced distributions. Specifically, we use $\alpha=0.9$ by default to simulate mild skewness, and decrease it to 0.3 to model more severe distribution shifts. During local training, each benign client performs 2 epochs with a learning rate of 0.03, while the global training proceeds for 1200 rounds. For conciseness, we omit the less critical parameter details and follow [17] for all settings not explicitly specified.
@@
Backdoor Attack Configurations. To ensure a comprehensive evaluation, we consider both single-attacker and multi-attacker scenarios, each covering various attack strategies and backdoor types. (1) In the single-attacker scenario, we adopt a range of attack baselines, including the naive Vanilla training [12], the stealthy PGD attack [37], and the latest state-of-the-art methods Neurotoxin [48] and Chameleon [8]. We adopt three diverse types of backdoor triggers, ranging from the traditional visible trigger (BadNet [12]), to the invisible trigger (Blend [7]), and the natural feature-based trigger (Semantic [2]), with increasing levels of imperceptibility. (2) In the multi-attacker scenario, we increase the ratio of attackers from 30% to 70%. To ensure comprehensive coverage, we include both aligned and divergent attack goals: Uniform [9] with homogeneous attackers, DBA [41] with a shared target but distributed triggers, and NBA [29] with non-cooperative attackers using distinct backdoors. Across all settings, the poisoning stage uses a learning rate of 0.03 by default, consistent with benign training, and reduces it by half for more challenging stealthy evaluations. The backdoor target is class 0 by default. Similar to Indicator, the malicious client launches the attack at the 1000th round, with the poisoning action lasting for 200 rounds.
@@
Backdoor Detection Configurations. We evaluate six representative federated backdoor detection methods that span a range of strategies, from classical model-space analysis to more advanced and proactive techniques: We begin with traditional methods that rely on the statistical properties of model updates: (1) MultiKrum [5] detects backdoor models based on the Euclidean distance among client updates. (2) Foolsgold [9] identifies malicious clients by analyzing pairwise cosine similarity to detect overly aligned update directions. (3) Rflbat [39] projects model updates into a PCA space and flags outliers as potential attackers. Then we further include more advanced approaches that incorporate prediction behavior or mitigation mechanisms: (4) Deepsight [34] evaluates similarity in model predictions and output neuron activations to reveal abnormal client behavior. (5) Flame [27] combines detection based on model similarity together with mitigation techniques such as gradient clipping and noise injection. Finally, we include the pioneering proactive method that shifts the detection paradigm: (6) BackdoorIndicator [17] injects OOD patterns into the global model and identifies backdoor models based on their reactions to these patterns.
@@
Configuration of Coward. Here we detail the key factors related to Coward. To construct the OOD planting set, we use EMNIST (expanded to three channels) when the main task is CIFAR-10 or CIFAR-100, and grayscale CIFAR-10 when the main task is EMNIST. The same planting set is used for Indicator for fairness. By default, the size of the planting set is 1000. For watermark injection, we use 20% of the planting set samples for trigger mapping, with a learning rate of 0.001 for 5 iterations. In our main experiments, we adopt WaNet [28] as the default watermark trigger. For detection, a fixed 5% threshold is used across all settings: clients with watermark accuracy below this are identified as malicious. The watermark target is 8 by default.
@@ 5.2 Main Results
We comprehensively evaluate all baseline detection methods from three aspects:
Resistance to FL Non-i.i.d. We evaluate all baseline methods under varying degrees of non-i.i.d. data across three datasets, using a vanilla backdoor setting with a fixed spot-pattern trigger and a learning rate matching that of benign clients. As shown in Table 1, Coward consistently achieves the best overall performance across all levels of heterogeneity, with TPRs above 95% and FPRs below 10.5% in all cases. Besides the overall conclusion, we draw several key observations: (1) Coward significantly reduces FPR across all settings compared to BackdoorIndicator, highlighting its advantage in mitigating OOD bias. A more detailed analysis is provided in Section 5.5. (2) Proactive methods significantly outperform passive ones under severe data heterogeneity. Taking the most challenging non-i.i.d. setting ($\alpha=0.3$) as an example, aside from our method, BackdoorIndicator also outperforms all passive baselines, achieving TPRs of 93% on CIFAR-10 and 96% on EMNIST. This result supports the robustness of proactive detection in heterogeneous environments. In contrast, passive methods exhibit significant performance degradation. For instance, the state-of-the-art passive method Flame sees its TPR drop from above 95% to below 10% on both CIFAR-100 and EMNIST, which aligns with our earlier analysis. (3) Some passive methods fail due to their unrealistic assumptions. For example, Rflbat performs poorly across all heterogeneity levels, likely because it relies only on the top two principal components of model parameters, losing critical information, especially under our challenging FL setup. Similarly, FoolsGold collapses on CIFAR-100, always judging all clients as benign (resulting in 0% TPR and 0% FPR). This failure stems from two factors: the high-dimensional 100-class prediction vector and its reliance on multi-attacker scenarios. These conditions lead to extremely low cosine similarity values, causing an all-as-benign judgment which is unhelpful for detecting malicious clients.
@@
Table 1: Defense results against vanilla single-client backdoor attack under different non-i.i.d. settings. Boldface indicates the best result value (excluding the collapsed case with TPR/FPR=0/0), while red highlights metric-specific failures based on a 50% threshold ($i.e.$, values above 50% for FPR and ASR, and below 50% for TPR and BA).
(a) Detection performance.
@@
Dataset NonIID MultiKrum FoolsGold Rflbat DeepSight Flame Indicator Ours TPR FPR TPR FPR TPR FPR TPR FPR TPR FPR TPR FPR TPR FPR CIFAR10 0.9 93.5 34.1 44.0 44.3 2.5 9.6 74.0 60.3 99.0 33.4 97.5 23.0 99.0 3.9 0.6 62.0 37.6 47.5 50.3 1.5 10.4 66.0 53.2 69.0 36.8 94.0 26.1 99.0 9.2 0.3 57.5 38.1 60.5 57.2 1.0 9.2 57.0 44.8 62.0 37.5 93.0 31.0 99.0 9.6 CIFAR100 0.9 98.0 22.4 0.0 0.0 12.0 22.6 39.0 32.6 97.0 22.6 95.0 47.0 98.0 2.9 0.6 96.0 22.7 0.0 0.0 11.0 22.4 55.0 47.8 95.0 22.8 81.0 47.8 100.0 4.3 0.3 2.0 33.1 0.0 0.0 6.0 23.3 75.0 65.3 5.0 32.8 76.0 43.6 98.0 8.6 EMNIST 0.9 100.0 22.2 52.0 39.1 16.0 22.7 57.0 45.9 100.0 22.0 99.0 57.8 98.0 1.6 0.6 98.0 22.4 35.0 41.7 11.0 23.1 56.0 49.4 97.0 22.3 96.0 52.7 98.0 1.6 0.3 9.0 32.3 40.0 48.2 1.0 25.1 49.0 47.0 9.0 32.1 96.0 45.6 98.0 8.3
(b) Attack success rate and benign accuracy.
Dataset NonIID MultiKrum FoolsGold Rflbat DeepSight Flame Indicator Ours ASR BA ASR BA ASR BA ASR BA ASR BA ASR BA ASR BA CIFAR10 0.9 34.1 92.1 91.3 92.1 91.3 91.8 58.2 91.9 9.9 92.1 9.4 91.2 10.3 90.5 0.6 88.9 91.9 89.5 91.4 91.3 92.0 66.4 92.0 87.2 91.8 35.3 91.5 9.9 90.8 0.3 90.2 91.7 84.5 90.8 91.2 91.9 72.7 91.8 88.5 91.8 24.2 91.5 10.0 90.4 CIFAR100 0.9 1.1 69.7 82.9 69.9 83.2 69.7 46.8 69.8 1.1 69.7 1.3 69.0 1.0 69.1 0.6 1.0 69.9 82.7 69.5 83.4 69.6 41.5 69.5 1.0 69.8 56.1 69.0 1.0 69.2 0.3 82.0 69.5 81.5 69.3 82.4 69.2 25.2 69.7 81.6 69.3 46.8 67.8 1.0 68.9 EMNIST 0.9 10.0 99.8 99.9 99.7 100.0 99.7 27.4 99.7 10.0 99.7 10.0 99.6 10.0 99.7 0.6 10.0 99.8 99.7 99.7 99.8 99.7 47.8 99.7 10.0 99.7 10.8 99.6 10.0 99.7 0.3 100.0 99.7 99.5 99.8 99.9 99.7 47.3 99.8 100.0 99.7 10.2 99.6 10.0 99.7
Resistance to Advanced Attacks. We evaluate all baselines under more challenging, advanced, and stealthy attack settings. Specifically, the malicious client adopts a more imperceptible noise-blended trigger and reduced local learning rate in all attack methods. We report detection performance in Table 2, while additional ASR and BA metrics are reported in Supplementary Table 5, as they exhibit trends consistent with those in Table 1. As the results show, our method achieves the best overall performance, with proactive detection showing strong robustness, while passive methods suffer substantial degradation. This highlights the advantage of the proactive paradigm in handling sophisticated attacks. (1) Among proactive methods, Coward consistently achieves the best performance, maintaining close to 100% TPR and sub-10% FPR across all advanced attack types, including state-of-the-art strategies such as Neurotoxin and Chameleon. BackdoorIndicator also performs strongly, with TPRs consistently exceeding 85% in most settings; however, it suffers from relatively high FPRs.(2) For passive methods, performance degradation is anticipated under such stealthy attacks, since the attacker introduces only subtle changes to the model, which reduces its suspiciousness. As an instance, MultiKrum, which selects representative updates based on Euclidean distance, becomes highly ineffective, with TPRs falling below 12.5% across all stealthy attack settings. Flame, which filters out updates with large angular deviations, also suffers notable degradation, with TPRs dropping below 30.5% on the CIFAR-10 dataset. In contrast, DeepSight shows better resilience, which performs consistently with TPRs over 40% in most attack cases. This probably stems from its joint use of both parameter and prediction discrepancies.
Table 2: Detection performance against stealthy single-client backdoor attacks with advanced training strategies. Result highlighting (in boldface and in red) follows the same rules as those described in Table 1. NTX and CML stand for Neurotoxin and Chameleon, respectively.
Dataset Attack MultiKrum FoolsGold Rflbat DeepSight Flame Indicator Ours TPR FPR TPR FPR TPR FPR TPR FPR TPR FPR TPR FPR TPR FPR CIFAR10 Vanilla 5.5 32.7 53.5 44.4 3.0 24.2 77.9 64.0 30.5 29.9 93.0 34.7 100.0 3.3 PGD 5.5 32.7 53.5 44.6 2.0 24.5 75.5 61.0 12.0 31.8 92.0 34.1 97.5 4.2 NTX 6.0 32.7 15.5 43.1 2.5 24.3 69.2 58.5 30.0 29.9 52.5 31.7 99.5 7.8 CML 8.0 32.4 46.5 44.1 3.5 23.6 39.0 36.5 10.0 32.2 87.0 33.8 94.5 9.1 CIFAR100 Vanilla 6.5 32.6 0.0 0.0 3.5 23.7 48.7 44.4 7.5 32.5 91.0 45.3 100.0 1.9 PGD 6.5 32.6 0.0 0.0 2.5 23.4 46.7 43.0 2.0 33.1 88.0 45.7 100.0 5.8 NTX 7.0 32.6 0.0 0.0 4.5 23.2 51.3 42.7 2.5 33.1 96.0 46.9 100.0 9.4 CML 0.0 33.3 0.0 0.0 5.0 24.3 41.0 34.7 0.0 33.3 98.5 47.5 98.2 7.3 EMNIST Vanilla 8.5 32.4 12.0 35.7 1.0 24.6 24.0 37.3 44.5 28.3 88.3 45.7 100.0 3.6 PGD 10.0 32.2 11.0 35.2 2.0 11.7 44.0 46.9 37.0 31.4 88.5 46.0 100.0 4.3 NTX 10.5 31.9 2.5 36.3 1.5 12.0 38.5 43.1 44.5 17.2 88.6 45.3 99.0 3.6 CML 12.5 31.9 21.5 37.2 3.0 24.8 37.5 42.7 98.0 22.3 91.5 51.7 99.5 2.4
Robustness to Multiple Attackers. Table 3 presents detection performance under multi-attacker scenarios. Additional ASR and BA metrics are reported in Supplementary Table 6, as they exhibit trends consistent with those in Table 1. As shown in this table, proactive methods exhibit better overall performance across all three settings. This is expected as the detection mechanism of proactive approaches is independent of the number of attackers. Among them, our method achieves the best overall performance, consistently maintaining TPRs above 99% and FPRs below 15%. BackdoorIndicator also performs well but becomes less effective when the attacker proportion reaches 70%. In such cases, a temporary detection failure allows the increased number of attackers to rapidly reinforce the global backdoor, resulting in a higher ASR. Moreover, the retained backdoor could be further maintained by the indicator’s maintenance effect. Among passive baselines, FoolsGold performs best under Uniform and DBA attacks, achieving a TPR of 100%, benefiting from its assumption of highly similar, coordinated malicious updates. However, it fails in uncoordinated scenarios, where attackers pursue different targets. In such cases, the TPR drops below 40% as the method’s core assumption of shared attack objectives no longer holds. Outlier-based methods such as MultiKrum and Flame become ineffective when attackers control 70% of clients, with TPRs dropping under 42% in all cases. These methods assume malicious updates are outliers, which fails when malicious clients form the majority.
Table 3: Detection performance under multi-client backdoor attacks on the CIFAR10 dataset. Result highlighting (in boldface and in red) follows the same rules in Table 1.
Strategy Atk% MultiKrum FoolsGold Rflbat DeepSight Flame Indicator Ours TPR FPR TPR FPR TPR FPR TPR FPR TPR FPR TPR FPR TPR FPR Uniform 30% 100.0 71.4 100.0 34.2 39.4 22.5 3.7 14.6 100.0 14.3 99.2 29.5 100.0 6.7 50% 3.0 97.0 100.0 27.1 31.3 0.0 1.1 24.0 21.5 57.3 99.1 16.3 100.0 8.2 70% 29.0 98.9 100.0 12.8 14.3 11.8 2.4 54.8 31.3 57.8 99.4 44.6 100.0 11.2 DBA 30% 100.0 57.4 100.0 32.9 28.0 15.4 5.6 12.4 100.0 14.3 98.2 18.4 100.0 9.7 50% 22.8 77.2 100.0 26.7 29.2 0.0 2.9 20.1 60.5 18.9 98.6 16.0 100.0 9.9 70% 31.1 94.1 100.0 12.8 9.5 2.8 3.4 44.7 42.0 34.8 98.8 15.8 100.0 11.2 NBA 30% 100.0 57.1 37.5 48.4 78.5 70.8 2.8 31.2 55.2 18.9 97.8 39.4 98.7 7.1 50% 63.7 36.3 33.3 54.6 16.5 15.8 1.4 48.1 19.2 40.6 93.7 38.2 98.5 15.0 70% 36.9 80.7 23.4 62.7 0.1 0.0 1.3 78.8 10.6 74.8 81.3 14.0 86.2 14.6
Table 4: Performance on different OOD planting sets. Result highlighting (in boldface and in red) follows the same rules in Table 1.
@@
5.3 Ablation Study
We investigate key factors that influence watermark planting and detection: (1) the choice of OOD planting set, (2) the pairing of watermark and backdoor triggers, and (3) the selection of detection thresholds. For a more comprehensive analysis, please refer to Supplementary Sections B and C.
Robustness to Choices of OOD Datasets. Table 4 presents detection performance across diverse OOD planting sets, including task-similar data, distinctly different data, and an extreme case with synthetic Gaussian noise. We observe that Coward consistently maintains low FPR across all settings, while Indicator’s performance degrades as the OOD dataset diverges from the main task. Using the most challenging main task, the CIFAR-100 dataset, as an example, Coward’s FPR remains below 6.3% across all planting sets. In contrast, Indicator’s FPR increases sharply from 44.6% on CIFAR-10 to 60% with Gaussian noise. This confirms our earlier analysis of Indicator’s increased vulnerability to OOD bias. Overall, these results show that Coward is more robust to OOD data selection, offering greater flexibility in real-world deployment.
Robustness to Choices of Triggers. Table 5 summarizes performance under various combinations of trigger types. The attacker’s trigger varies in increasing stealthiness: from a visible pixel pattern, to an invisible blend pattern, and finally to an implicit semantic pattern. The watermark trigger also spans a range of types, including four visible patterns (diagonal, square, triangular, and noise), as well as two invisible variants: a mosaic pattern and the sample-specific WaNet pattern. We see that Coward performs consistently well across all settings, achieving a TPR above 96.5% and an FPR below 17.3%. These results demonstrate our method’s robustness to diverse triggers and the generalizability of the multi-backdoor collision effect.
Table 5: Detection performance with different combinations of watermark trigger (S) and attackers’ client-side malicious trigger (A) on the CIFAR-10 dataset.